OpenStudy (anonymous):

Find the exact value of the expression sec[sin^-1((-sqrt2)/2)]

OpenStudy (anonymous):

\[\sec [\sin^{-1} ((-\sqrt{2})/2)]\]

OpenStudy (anonymous):

Is it "no solution" because you can not determine the x value

OpenStudy (anonymous):

and/or because you can not determine which quadrant the expression lies.

OpenStudy (anonymous):

or is the answer \[\sqrt{2}\]